OpenStudy (anonymous):

find the domain of the function. f(x) = -2x

OpenStudy (anonymous):

All real numbers.

OpenStudy (anonymous):

how do you write that.... bracket-2,infinity parentheses?

OpenStudy (anonymous):

It would be (-infinity,infinity)

OpenStudy (anonymous):

The domain is the input of the function. You can put anything in for x in this case.

OpenStudy (anonymous):

Contrast this with 1/x. That would be all real numbers other than zero since you can't divide by zero.

OpenStudy (anonymous):

can you help me with one more hard one

OpenStudy (anonymous):

find f(-5)= f(0)= f(0.5= f(1)= f(4)=

OpenStudy (anonymous):

You need to figure out which part of the function to use. f(-5) means x = -5, -5 < 0 so we use -2x; -2(-5) = 10 f(0) means x = 0, x =0 so we use x - 4, 0 - 4 = -4 The same reasoning applies to the next three.